OpenStudy (anonymous):

1+a/5= -1 (One plus a over 5 = negative 1 )

OpenStudy (katrinakaif):

Do you need to solve for a?

OpenStudy (anonymous):

Its a 2 step equation. And yes.

OpenStudy (katrinakaif):

Well if you want to solve for "A" it would be = to-10.. Why? Because 1 + a/5 = -1 Bring the +1 to the other side by substracting it.. 1 + a/5 = -1 a/5 = -1 - 1 a/5 = -2 Now mulitply by the recirpocal of 1/5 which is 5/1 on the other side.. a/5 = -2 a = -2x5/1 a = -10 I hope this is what you wanted.. O_o

OpenStudy (anonymous):

Thank you!